OpenStudy (anonymous):

I need to understand how to solve (x^6)^1/3 @dumbsearch2 can you help?

OpenStudy (dumbsearch2):

I'll try :)

OpenStudy (anonymous):

Thank You. I am new to openstudy. I was told about so I thought I would give it a try.

OpenStudy (anonymous):

\[(x^n)^m = x^{n\cdot m}\]

OpenStudy (anonymous):

@pgpilot326 okay can you do that with numbers it doesn't have to be with the same numbers i gave. can you make the problem different from what i gave?

OpenStudy (anonymous):

sure... for example:\[(x^2)^3=(x^2)\cdot (x^2)\cdot (x^2)=x^6 = x^{2\cdot 3}\]

OpenStudy (anonymous):

okay okay I get it now but it doesn't matter if it's a fraction outside of the parenthesis right?

OpenStudy (anonymous):

@pgpilot326

OpenStudy (anonymous):

nope, doens't matter what kind of number is inside or outside. the property holds regardless. you can even switch ... (so long as the operation is commutative) \[(x^n)^m = (x^m)^n = x^{n\cdot m}= x^{m\cdot n}\]

OpenStudy (anonymous):

so did you work it out?

OpenStudy (anonymous):

Yes I did. You are a great help. I wish i had something to give you for your help. Thank you Thank You! @pgpilot326

OpenStudy (anonymous):

your appreciation and learning is more than enough!!! keep up the good work!!!

OpenStudy (anonymous):

Thank You @pgpilot326

OpenStudy (anonymous):

@Nikia_perez You are so welcome!